Calculate the amount of agarose and 10,000x SYBR safe needed to make 15 mL of 1% agarose 1x SYBR safe  

Amounts for 15ml 1% agarose

1x TAE

15mL

Agarose powder (g)

10,000x SYBR safe
Answer #1

1% agarose = 1g/100mL solution .

Thus for 15mL solution we need = (1g/100mL)×15mL = 0.15g of agarose powder.

For SYBR safe :

C1V1 = C1V2

10000X × V1 = 1X × 15mL

V1 = (1×15)/10000 = 0.0015 mL

Thus we require, 0.0015 mL of 10000X SYBR safe

And 0.15g of agarose powder for the required TAE .